OpenStudy (jazzyfa30):

cos-105 A) (squareroot6)+(squareroot2)/4 B)(squareroot6)-(squareroot2)/4 C)(squareroot2)-(squareroot6)/4 D)-(squareroot6)-(squareroot2)/4

OpenStudy (jazzyfa30):

@ganeshie8

ganeshie8 (ganeshie8):

hint : \(\cos (-\theta) = \cos (\theta)\)

ganeshie8 (ganeshie8):

hint2 : \(105 = 60 + 45\)

OpenStudy (jazzyfa30):

I use cosAcosB-sinAsinB or CosAcosB-sinAsinB

OpenStudy (jazzyfa30):

^^^^^^^^^^^^+

ganeshie8 (ganeshie8):

\(\cos (A +B) = \cos A \cos B \color{red}{-} \sin A \sin B\)

ganeshie8 (ganeshie8):

\(\cos (-105) = \cos (105) = \cos(60+45) ~= ~?\)

OpenStudy (jazzyfa30):

cos60cos45-sin60sin45

ganeshie8 (ganeshie8):

Yep ! look up ur unit circle

OpenStudy (jazzyfa30):

\[1/2*\sqrt{2}/2-\sqrt{3}/2*\sqrt{2}/2\]

ganeshie8 (ganeshie8):

Perfect ! simplify...

OpenStudy (jazzyfa30):

\[\sqrt{2}/4-\sqrt{6}/4\]

OpenStudy (jazzyfa30):

so my answer is c

ganeshie8 (ganeshie8):

\(\large \color{Red}{\checkmark }\)

ganeshie8 (ganeshie8):

good job !!